OpenStudy (jessiegonzales):

MEDAL!!!!!! 10. Solve the equation or formula for the indicated variable. a=21^3w, for w A. w=a/21^3 B. w=a/21 C. w=1^3/2a D. w=a/3-2

OpenStudy (jessiegonzales):

@zepdrix

OpenStudy (jessiegonzales):

@nincompoop

OpenStudy (zale101):

\(\Large a=21^3w\) like this?

OpenStudy (jessiegonzales):

yes

OpenStudy (kropot72):

Or is it this? \[\large a=21^{3w}\]

OpenStudy (jessiegonzales):

no

OpenStudy (jessiegonzales):

\[a=21^3 w\]

OpenStudy (kropot72):

Divide both sides by 21^3, to solve for w.

OpenStudy (jessiegonzales):

a=21^3w/ 21/3??

OpenStudy (nurali):

a=21^3w Divide both side by 21^3 a/21^3=21^3/21^3w w=a/21^3